OpenStudy (anonymous):

Determine whether the function is one-to-one. f(x)= x^(4)+5, 0<=x<=2

OpenStudy (anonymous):

one to one function

OpenStudy (anonymous):

most polynomials are

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

in order to solve this put 0 1 and 2 into the equation and solve if you get different y values then it is one to one

OpenStudy (anonymous):

you need to show f(a)=f(b) => a=b, so put a^4 +5 = b^4+5 => a^4 = b^4, since 0<=x<=2 => a=b, so your function is 1-1. Note however, if your domain did not force x to be between 0 and 2, this function would not be 1-1.